OpenStudy (anonymous):

The attendance on the first day of a carnival was 425 people. The attendance on the third day was 575 people. Assuming attendance will increase linearly each day, how many people will attend the carnival on the sixth day? Question 3 answers 650 725 800 875

OpenStudy (amistre64):

1,425 -(2,575) -------- -1,-150; slope = y/x = 150 y=150(x-1)+425

OpenStudy (amistre64):

725

OpenStudy (anonymous):

thanks (: <3